# PointRecon: Online Point-based 3D Reconstruction via Ray-based 2D-3D Matching

### v1

@@ 4.2 Experiment Results
ScanNet. ScanNetv2 (Dai et al., 2017) is an indoor RGB-D video dataset comprising scans of 1,613 rooms, with 1,201 rooms designated for training, 312 for validation, and 100 for testing. We compare our results with previous work in terms of both 3D mesh quality (Table 1) and the 2D depth map quality (Table 2). We generate two versions of meshes with TSDF fusion resolutions set at 2cm and 4cm, respectively. As illustrated in the tables, our model achieves a high overall score (measured by Chamfer distance and F-score), ranking at the top among online methods. While our recall score is notably high, our precision is relatively low, indicating the presence of noisy and redundant surfaces in the reconstructed mesh. This issue partly arises from our method not incorporating any post-hoc smoothing steps. Additionally, our heuristic in the merging step, which retains points from older camera views, may also contribute to the increased noise levels in the scene. Future work will focus on developing improved smoothing techniques.
[…]
Offline COLMAP (Schönberger et al., 2016) / 0.069 0.135 0.102 0.634 0.505 0.558
Atlas (Murez et al., 2020) / 0.084 0.102 0.093 0.598 0.565 0.578
VoRTX (Gao et al., 2023) / 0.082 0.062 0.072 0.688 0.607 0.644
CVRecon (Feng et al., 2023) / 0.077 0.045 0.061 0.753 0.639 0.690
Online DeepVMVS (Duzceker et al., 2021) 37 0.076 0.117 0.097 0.451 0.558 0.496
NeuralRecon (Sun et al., 2021) 90 0.128 0.054 0.091 0.684 0.479 0.562
TF (Bozic et al., 2021) 326 0.099 0.078 0.089 0.648 0.547 0.591
SimpleRecon (Sayed et al., 2022) 72 0.078 0.065 0.072 0.641 0.581 0.608
VisFusion (Gao et al., 2023) 90 0.105 0.055 0.080 0.695 0.527 0.598
PointRecon (4cm) 618 0.059 0.078 0.068 0.576 0.645 0.607
PointRecon (2cm) 618 0.056 0.073 0.065 0.599 0.675 0.633
Table 1: Mesh Evaluation. Mesh reconstruction quality for the ScanNetv2 test split, following the evaluation protocol of Atlas (Murez et al., 2020). Offline methods assume access to the entire sequence and do not support local updates. The numbers in brackets indicate the TSDF Fusion resolution. While we present results for both 2cm and 4cm resolutions (with 4cm commonly used in volumetric methods), our point-based representation is not constrained by any specific TSDF Fusion resolution.
[…]
Offline COLMAP 0.264 0.137 0.138 - 83.4 87.1
Atlas 0.123 0.065 0.045 - 93.6 99.9
VoRTX 0.092 0.058 0.036 - 93.8 95.0
CVRecon 0.078 0.047 0.028 - 96.3 -
FineRecon 0.069 0.042 0.026 86.6 97.1 97.2
Online NeuralRecon 0.106 0.065 0.031 - 94.8 90.9
TF 0.099 0.065 0.042 - 93.4 90.5
SimpleRecon 0.083 0.046 0.022 - 95.4 94.4
PointRecon (4cm) 0.085 0.054 0.022 71.9 96.4 94.8
PointRecon (2cm) 0.087 0.055 0.024 72.1 96.2 94.6
Table 2: Depth Map Evaluation. Depth map quality for the ScanNetv2 test split, rendered from the reconstructed 3D meshes. We follow the evaluation protocol of Atlas (Murez et al., 2020). - indicates that the metric was not reported in the original paper. The number in parentheses represents the TSDF Fusion resolution.
Profiling. We profile the inference speed of our model’s major components on a single A100 GPU. On average, the image encoder contributes a negligible 1.8ms per image; the monocular depth predictor takes 8ms; the scene update step takes 215ms; the depth prediction with feature matching requires 276ms; and the point merging takes 118ms. The most computationally expensive part is the transformer U-Net, used in both the scene update and depth prediction steps. The merging step slows down as the point cloud grows. In future work, we aim to simplify the network to improve its speed.
Ablation. We analyze the contributions of different components of our model by evaluating the quality of the depth maps produced by the components on the ScanNet test split key frames. To assess the necessity of the scene update step, we train an additional model without this component. As shown in Table 3, feature matching significantly improves depth prediction accuracy compared to monocular prediction. Point merging further enhances the accuracy by leveraging predictions from multiple cameras, mitigating ambiguities that a single camera might face. The confidence-based merging results in a higher-quality point cloud. The scene update step is also crucial, as it refines the point cloud, leading to more accurate feature matching and merging in subsequent steps.
@@
(a) NeuroRecon
(b) SimpleRecon
(c) Ours
(d) Ground truth
(e) NeuroRecon
(f) SimpleRecon
(g) Ours
(h) Ground truth
Figure 4: Visualization of generated meshes. Our method produces more accurate reconstructions compared to NeuroRecon, but the meshes appear slightly less smooth due to the absence of explicit smoothing constraints in our approach. More reconstruction results can be viewed on our project page: https://arthurhero.github.io/projects/pointrecon/
@@ 5 Conclusion
We propose PointRecon, an online, point-based 3D reconstruction method that incrementally updates a global point cloud from posed monocular RGB images. Our method enables flexible local updates without requiring predefined resolution or scene size, while ensuring surface consistency across multiple views. Central to our approach is the novel ray-based 2D-3D matching strategy, which improves the robustness of feature matching, achieving state-of-the-art performance. However, due to the lack of post-hoc smoothing and heuristics, the reconstructions can still exhibit noise in certain areas. In the future, we aim to improve surface quality by incorporating advanced denoising techniques to make the results cleaner and more refined.

### v2

@@ 4.2 Experiment Results
ScanNet. ScanNetv2 [7] is an indoor RGB-D video dataset comprising scans of 1,613 rooms, with 1,201 rooms designated for training, 312 for validation, and 100 for testing. Each scan typically includes thousands of image frames, with approximately 1/10 of the frames designated as keyframes. We compare our results with previous work in terms of both 2D depth map quality (Table 1) and 3D mesh quality (Table 2), where the depth maps are rendered from the reconstructed 3D meshes. To showcase the flexibility of our point representation, we generate meshes at two TSDF fusion resolutions, 2cm and 4cm, demonstrating that our approach is not constrained by TSDF resolution. As shown in the tables, our model achieves competitive performance among online methods, particularly excelling in depth map metrics such as “Abs Diff” and “$\delta<1.25$,” and delivering comparable results in Chamfer distance and F-score. Additionally, our method uniquely offers resolution flexibility and viewpoint consistency, attributes not captured by standard evaluation metrics. A notable strength of our approach is its top-ranking recall score among all online methods, highlighting our model’s ability to achieve high coverage of ground-truth surfaces. However, a lower precision score reflects some surface redundancy in the reconstructed meshes. This issue stems from the current merging heuristic, which retains points from earlier camera views. While this strategy ensures robust coverage, it may also preserve outdated or less accurate points, leading to redundant surfaces. Addressing this limitation is a promising direction for future optimization.
Profiling. Following [26], we measure the inference speed of our model on a single A100 GPU. On average, the image encoder incurs a latency of 1.8ms per image. The monocular depth predictor requires 8ms, the scene update step takes 215ms, depth prediction with feature matching takes 276ms, and the point merging step adds 118ms. The most computationally intensive component is the transformer U-Net, used in both the scene update and depth prediction steps. The merging process also becomes slower as the point cloud grows. While point-based representations eliminate grid constraints, they introduce computational challenges, particularly in $K$ nearest neighbor searches, which become slower as the point cloud scales. To enhance efficiency, future work will prioritize simplifying the network architecture and further reducing the point cloud size. It is important to note that, despite these computational demands, our approach remains an online method capable of continuously updating the scene with new inputs. In contrast, offline methods require access to the entire sequence before reconstruction, learning features from all images jointly, which inherently imposes an upper limit on sequence length.
@@
Ablation. We analyze the contributions of different components of our model by evaluating the quality of the depth maps produced by these components on the ScanNet test split key frames. To assess the necessity of the scene update step, we train an additional model without this component. As shown in Table 3, feature matching significantly improves depth prediction accuracy compared to monocular prediction. Point merging further enhances the accuracy by leveraging predictions from multiple cameras, mitigating ambiguities that a single camera might face. The confidence-based merging results in a higher-quality point cloud. The scene update step is also crucial, as it refines the point cloud, leading to more accurate feature matching and merging in subsequent steps.
@@ 5 Conclusion
We propose PointRecon, an online, point-based 3D reconstruction method that incrementally updates a global point cloud from posed monocular RGB images. Our approach offers flexible local updates without the need for predefined resolutions or scene sizes, while ensuring surface consistency across views. A key contribution is the ray-based 2D-3D matching strategy, which improves feature matching robustness. Future work will focus on simplifying the architecture, reducing point cloud size, and refining the merging algorithm to enhance efficiency and performance.